If a pharmacist needs 70L of a 50% alchohol solution but only has a 30% alchohol solution and a 80% alchohol solution how many L of each would make a 50% solution?
Answer by stanbon(75887)

If a pharmacist needs 70L of a 50% alchohol solution but only has a 30% alchohol solution and a 80% alchohol solution how many L of each would make a 50% solution?
--------------------------------------
Equation:
alcohol + alcohol = alcohol
0.30x + 0.80(70-x) = 0.50*70
Multiply thru by 100 to get:
30x + 80*70 - 80x = 50*70
-50x = -30*70
x = 42 L (amt. of 30% solution needed in the mixture)
---
70-42 = 28 L (amt. of 80% solution needed in the mixture)
============================================================
Cheers,
Stan H.